OpenStudy (anonymous):

What is the value of the expression when a = 3 and b = 2? 3a^3 + 2b^3

OpenStudy (anonymous):

3(3)^2 + 2(2)^2

OpenStudy (anonymous):

Okay, so this is what I got. 3*3*3 = 27 27 * 3 = 81 2*2*2 = 8 8*2 = 16 81+ 16 = 97

OpenStudy (anonymous):

It should look like this

OpenStudy (anonymous):

Is that right? ^^^

OpenStudy (camerondoherty):

Well do what we did last time: a = 3 b = 2 So we substitute a for 3 and b for 2 \[3(3)^3 + 2(2)^3\] So Whats 3*3*3? Whats 2*2*2?

OpenStudy (anonymous):

Scroll up I did it, can you check it for me? :)

OpenStudy (camerondoherty):

Yep its correct! Good Job!

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Yay!! Omg you helped me so much!! xD Now I understand it! c: Thanks so much, both of you :D

OpenStudy (camerondoherty):

You're Welcome c; Though you did much of the work

OpenStudy (anonymous):

I passed my test!! :D You really helped me understand this

OpenStudy (anonymous):

No problem anytime

OpenStudy (anonymous):

Great

OpenStudy (camerondoherty):

Tag me i any of your questions c: I'm always here to help!